ANALYSIS OF THE QUASI-LINEARITY OF THE B̃ ^{1}A_{1} STATE OF SiH_{2} and SiD_{2} RADICALS

Y. MURAMOTO, \underline{H. ISHIKAWA} and N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578, Japan.

Recently we have observed OODR transitions terminating to the B̃ ^{1}A_{1} state of SiH_{2} and SiD_{2}. The OODR spectra observed indicate the quasi-linear behavior in the B̃ state based on the following features. Even or odd-v_{2} levels exclusively appear in the OODR via an intermediate K_{a}=odd or even rotational level. This even/odd-v_{2} progression is a typical pattern in the case of bent-linear transition. That is, SiH_{2} acts as a linear molecule in the B̃ state and its bending vibration is doubly degenerate. Thus there exist a vibrational angular momentum ℓ. The intensity pattern comes from a selection rule of a c-type transition; Δ(ℓ − K_{a}) = ±1. Bending excited levels exhibit negative g_{22}-values. This is an indication of the double minimum potential. If an electronic state of interest was doubly degenerate, there would be a possibility of Renner-Teller effect. However, the B̃ state correlates to ^{1}Σ^{−} state in the linear configuration. Thus, this double minimum potential originates from the quasi-linearity. The height of the barrier to the linearity was calculated to be about 200 cm^{−1}.^{a} It was confirmed that there are only ℓ(K_{a}) = 0 rotational levels in the v_{2} = 0 level. This means that the v_{2} = 0 level is located above the barrier to linearity. It is very probable considering the very low barrier height. In addition, the absolute value of the g_{22} constant rapidly decreases as the v_{2}-value increases. These observations strongly support the quasi-linearity of the B̃ state. In the presentation, results of the vibrational analysis on the quasi-linearity will be discussed.

aI. Tokue, private communication